OpenStudy (anonymous):

how would you factor 2r^3+250?

OpenStudy (anonymous):

\[2(r^3+125)=2(r^3+5^3)=2(?\] \[a^3+b^3=\left( a+b \right)\left( a^2-ab+b^2 \right)\]

OpenStudy (anonymous):

From Mathematica:\[2 r^3+250=2 (r+5) \left(r^2-5 r+25\right) \]